V = 1/3 Pi r²h = 1/3Pi * 10² *24 = 240 Pi
V = 753.98 m³
SA = Pi r² + Pi r√2r + h²
SA = Pi (10² + 10√(10^2 + 24^2 )) = Pi (100 + 10√676)
SA =

How would I get the rest of this or did I even do it right?

Answer by stanbon(75887)

V = 1/3 Pi r²h = 1/3Pi * 10² *24 = 8*314 = 2512
V = 753.98 m³
SA = Pi r² + Pi r√(r^2 + h²)
SA = Pi (10² + 10√(10^2 + 24^2 )) = Pi (100 + 10√676)
SA = Pi(360)
================
Cheers,
Stan H.